HUBBLE SPACE TELESCOPE DAILY REPORT #1156 PERIOD COVERED: 0700 EST Daily Status Report as of 160/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 5570 (Cycle 4 Calibration: Earthflats, Large Set) The WF/PC-2 was used to make a set of Earth flat calibration images using different filters in order to construct high quality flat fields of the OTA illumination pattern. A total of eight images were received and displayed. Three were reported to be underexposed, and the others were okay. 1.2 Completed WF/PC-2 5476 (Continuum Snapshot of 3CR Radio Galaxies) The WF/PC-2 was used to make snapshot images of radio galaxy 3C207.0 as part of a continuum survey and to study the galaxy's morphology. All of the images were received and displayed. They were reported to look okay. 1.3 Completed FOS/BL 5500 (Energy Balance and Emission Lines in DQ Herculis Type Stars: Cycle 4 Medium) The Faint Object Spectrograph (Blue Side) was used to make ultraviolet spectral observations of the eruptive binary star YY Draconis. The observations executed, and was stored on the Science Tape Recorder #1. After the tape playback, it was reported that seven observations were received, and they looked okay. More data still remains on Science tape Recorder #3. (See Operations Note 2.1) 1.4 Completed WF/PC-2 5369 (HST Medium-Deep Survey: Cycle 4 Medium, High Galactic Latitude) The WF/PC-2 made a series of observations of high latitude objects as part of a medium deep sky survey. The observations were taken in parallel with other prime science observations. The first four observations were played back from Science Tape Recorder, and they were reported to look okay. More data still remains on Science Tape Recorder #3, and will be analyzed later. (See Operations Note 2.1) 1.5 Completed WF/PC-2 5655 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 2 Flats Part 2) The WF/PC-2 was used to make ultraviolet flats. Two images were received and displayed. No anomalies were noted. 1.6 Completed WF/PC-2 5446 (The Nuclei of Nearby S0 and Spiral Galaxies) The WF/PC-2 was used to make observations of galaxy NGC-3184. The observations were executed as scheduled, and the science data remains of Science Tape Recorder #3. It will be analyzed later. No anomalies were noted. 1.7 Completed FOS/BL 5438 (Io's Atmosphere: UV Spectroscopy of Neutral Components) The Faint Object Spectrograph (Blue Side) was used to make far ultraviolet spectrographic observations of Jupiter's moon Io. The first observations was received in real-time. The OSS computed and successfully uplinked a corrective pointing maneuver. No anomalies were noted. 1.8 Completed WF/PC-2 5397 (Determination of the Extragalactic Distance Scale: Cycle 4 High) The WF/PC-2 was used to make images of galaxy NGC-4535 to be used in measuring the periods of Cepheid variable stars within the galaxy. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 10 Successful : 10 At 160/0251 the reacquisition attained fine lock. However, the star presence was lost near the end of a sequential update. The Take Data Flag (TDF) did not go up, and the shutter did not open for subsequent FOS observations. 2.2 FHST Updates: Scheduled :12 Successful :12 2.3 Significant HSTARs: None 2.4 Operations Notes: There were no outstanding problems onboard the HST during the previous operational day. 3.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-